An amount invested at 10% compounded semiannually will double in ten years. I believe this problem is true. Am I right?
Answer by stanbon(48551)

A=P(1+r/n)^(nt); This is the general compounding formula.
If the investment amount(P) doubles, A will equal 2P
2P=P(1+(0.10/2))^(02(t))
2=(1.05)^(2t)
Take the ln of both sides to get
ln(2)=(2t)[ln(1.05)]
2t=14.21
t=7.105years
Cheers,
Stan H.